# LocMoE+: Enhanced Router with Token Feature Awareness for Efficient LLM Pre-Training

### v1

@@ 3.3.2 Theoretical Explanation
Hence, the traning of an input in the current step is valid if the class-discriminative patterns is correctly dispatched. We compute training success rate of input (i.e., the probability of the tokens in an input is correctly dispatched to experts) to measure the difference between TCR and ECR,
We follow the setup in [39] that the input ${\bm{x}}\in{\mathbb{R}}^{sd}$ with $s$ tokens is comprised of one class-discriminative pattern ${\bm{o}}_{1},\dots,{\bm{o}}_{n}\in{\mathbb{R}}^{d}$, with each decides the label $y\in[n]$, and $s-1$ class-irrelevant patterns ${\bm{r}}\sim{\mathcal{N}}$ (the distribution will be clarified below). To show the quantitative comparison of TCR and ECR in training success rate, we need following asssumptions of token patterns.
Assumption 3.1 (class-discriminative).
@@
${\mathbb{P}}({\bm{o}}_{i}\text{ sent to expert }i)=p_{i}\geq 1/n,\forall i\in[n].$
Assumption 3.2 (class-irrelevant).
@@
${\mathbb{P}}({\bm{r}}\sim{\mathcal{N}},{\bm{r}}\text{ sent to expert }i)=1/n,\forall i\in[n].$ (5)
@@
${\mathbb{P}}({\bm{r}}\sim{\mathcal{N}},\text{ the score of }{\bm{r}}\text{ is larger than }{\bm{o}}_{i}\text{ for expert }i)=q_{i},\forall i\in[n],$ (6)
@@
Assumption 3.1 assumes the valid token is uniformly distributed in training samples due to the massive amounts of data nowadays. Assumption 3.2 assumes the invalid tokens can be uniformly dispatched to experts since the invalid tokens do not provide supervised signal to router and experts in training. We consider such uniform settings are common assumptions in theoretical analyis.
Now we compute the training success rate of TCR and ECR.
Theorem 3.3.
Under Assumptions 3.1 and 3.2, the training success rate of TCR in each sample ${\bm{x}}$ is
@@
and the training success rate of ECR in each sample ${\bm{x}}$ is
@@
Remark 3.4.
In practice, we have constant number of experts [7], i.e., $n=\Theta(1)$, and $C<s$ to save computation cost. We explain the benefit of swicthing TCR to ECR during training based on Theorem 3.3 and feature distrution during training.
At the beginning of training, the model seldom learn the task. Then the feature of class-irrelevant tokens is nearly isotropy, e.g., uniformly distrbute around the sphere (see Appendix B), leading to $q_{i}=\Theta(1)$. The succed rate of TCR with the form $C/s$ is better than ECR with the form $e^{-s}$. Thus we should choose TCR with a large capacity $C=\Theta(s)$ to improve the success rate of training samples.
After training for some iterations, the experts can roughly distinguish the class-irrelevant and discriminative patterns, leading to $q_{i}\ll 1$ or $sq_{i}\leq C^{*}$ for some $C^{*}>0$ (see Appendix B). Then ECR with success rate nearly $1$ is better than TCR with the form $C/s$ as long as $C\geq 2C^{*}$. Thus we should choose ECR with a small capacity $C=\Theta(1)$ to improve the success rate of training samples.
Indeed, we find that Chowdhury et al. [39, the definition of $\ell^{*}$] consider the ECR setting and verify the benefit in sample complexity. They assume the maximum number of class-irrelevant patches that are close to class-discriminative patches are bounded, which has similar effect as $C^{*}$ in our scene.
3.4 Communication Optimization
4 Experiments
4.1 Experimental Setup
The details of experimental setup including datasets, environment, configuration, and metrics, can be seen in Appendix C.
4.2 Efficiency Promotion and Memory Footprint Reduction
As detailed in Section 3, we consistently use Top-1 routing to ensure the routing implementation aligns with our theoretical framework. The Baseline model utilizes a limited expert capacity mode instead of the groupedGEMM scheme, which avoids token dropping, with the capacity factor set to 1.1. LocMoE considers data distribution uniformity and estimates expert capacity using a lower bound formula derived from its theoretical conclusions in the first batch, maintaining it as a constant during subsequent training. LocMoE+ fixes the range of score sums, processes hidden states, and calculates current expert capacity. The subsequent analysis addresses the training time, convergence, and memory usage efficiency of these schemes on multiple sizes of Ascend clusters.
@@ 4.2 Efficiency Promotion and Memory Footprint Reduction
Figure 2 illustrates the time consumption of these methods during the first 1000 iterations of training. Due to initialization and some unstable factors, time consumption is recorded starting from the 5th iteration. The Baseline model’s time consumption is relatively stable. As iterations increase, LocMoE’s time consumption slightly decreases, particularly in 32N and 64N, consistent with the conclusion that locality loss is effective only when the number of experts is greater than or equal to the number of nodes. LocMoE+ incurs slightly higher time consumption than LocMoE due to the computational overhead of token rearrangement. However, as token features converge, the required tokens gradually decrease and stabilize, leading to a decline in time consumption, which remains stable in subsequent training processes. Overall, LocMoE+ reduces training time by 2.9% to 13.3% compared to LocMoE, and by 5.4% to 46.6% compared to the Baseline.
@@
We select 10 iterations at equal intervals from the training iterations to collect data on the time consumption of computation, communication, overlap, and idle periods, as shown in Figure 3(a). It is important to note that the data collection operation also introduces some overhead. After integrating LocMoE and LocMoE+, the time consumption of each component decreases, with a significantly greater reduction in computation overhead compared to communication overhead. Additionally, as the cluster size increases, the proportion of computation/communication overlap decreases, and the magnitude of the reduction in computation overhead diminishes. Figure 3(b) illustrates perplexity as a measure of convergence. The convergence curves of these approaches indicate normal loss convergence, with LocMoE+ not adversely impacting convergence.
@@
The proportional time consumption at the operator level is depicted in Figure 4. Among the components, AI CORE efficiently executes matrix multiplications and convolutions in AI algorithms; AI VECTOR CORE accelerates vector operations through parallel processing; MIX AIC integrates different types of operators and optimizes for multiple tasks; AI CPU is optimized in hardware and instruction sets to better support AI algorithms. LocMoE+ selects fewer tokens, resulting in a 17$\times$ performance improvement in the FFN MatMul operator compared to the Baseline and a 2.6$\times$ improvement compared to LocMoE. This leads to an overall 2.8$\times$ reduction in the cumulative time consumption of the MatMul operator and a 2.6$\times$ decrease in Cube computing load. However, the proportions of TopK and IndexPutV2, related to rearrangement, show a slight increase.
@@
Figure 5: The footprint recorded in one acquisition cycle with different schemes and cluster sizes.
We select a single iteration during the stable training period and describe the per-device memory usage (Allocated) using the first 100,000 samples from its memory monitoring, as shown in Figure 5. Overall, LocMoE+ achieves memory usage reduction of 4.57% to 16.27% compared to the Baseline and 2.86% to 10.5% compared to LocMoE. As cluster size increases, the proportion of computational overhead decreases, and the gap in memory usage narrows. Additionally, instantaneous memory peaks gradually disappear, and the fluctuation amplitude of short-term memory also diminishes.
4.3 The Performance of Downstream Tasks
To enhance the model’s conversational capabilities and adaptability to downstream task, we fine-tuned the pre-trained models. As shown in Figure 6, with sufficient supervised fine-tuning (SFT), LocMoE+ achieves an average improvement of approximately 20.1% in 16 sub-capabilities of Domain Task Capability, which is a portion of General and Domain-specific Assessment Dataset (GDAD), compared to the Baseline, and an increase of about 3.5% compared to LocMoE. The Rewriting and Summary capabilities show the highest improvement, with a 28.2% increase compared to the Baseline and a 6.7% increase compared to LocMoE. In the 13 tests of Domain Competency Exam, LocMoE+ demonstrates an average improvement of 16% relative to the Baseline and an average increase of approximately 4.8% compared to LocMoE. The IP Training in the digital communications domain shows the most significant improvement, with a 27.3% increase compared to the Baseline and a 3.0% increase compared to LocMoE. Among the 18 sub-capabilities of General Ability, LocMoE+ exhibits an improvement of about 13.9% relative to the Baseline and an average increase of 4.8% compared to LocMoE. The capability of Planning demonstrates the highest improvement, with a 26.8% increase compared to the Baseline and a 2.92% increase compared to LocMoE.
@@ 4.3 The Performance of Downstream Tasks
Table 1 presents the holistic evaluation results for multiple datasets, where GDAD-1 represents Domain Task Capability, and the other metrics follow accordingly. Notably, due to the 6:4 ratio of Chinese to English data in our incremental pre-training domain data and the 7:3 ratio in the fine-tuning data, LocMoE+ achieves an improvement of approximately 10.7% compared to the Baseline and 1.2% compared to LocMoE in the C-Eval [40] evaluation, despite the limited data available for training. During incremental training and fine-tuning, we incorporated substantial telecommunications domain knowledge, questions, and case studies. TeleQnA [41], the first benchmark dataset designed to evaluate the knowledge of LLMs in telecommunications, effectively measures the model’s capabilities in this domain. Consequently, LocMoE+ comprehensively surpasses both the Baseline and LocMoE on this specific dataset.
Table 1: Performance promotion obtained by LocMoE+ on different datasets.
@@
5 Conclusion
In this paper, we propose a novel MoE structure incorporating token feature awareness mechanism to enhance the training efficiency and performance of LLMs. Building upon the feature extraction component and locality loss of LocMoE, LocMoE+ introduces a routing strategy that combines TCR and ECR, theoretically demonstrating improved training success rates. By defining the affinity between tokens and experts, expert capacity can be dynamically reduced and further minimized. Model training experiments are conducted on Ascend clusters, leveraging the communication optimization characteristics of CoC. LocMoE+ achieves performance improvements up to 46.6% (32N) compared to the Baseline and 13.3% (32N) compared to LocMoE, while reducing memory usage by up to 16.27% and 10.5%, respectively. To evaluate model performance, all pre-trained models are fine-tuned using a mixture of domain-specific and general data, and assessed with the open-source datasets C-Eval and TeleQnA, and closed domain benchmark GDAD. In downstream tasks, LocMoE+ outperforms the Baseline by 14.1%, 10.7%, and 9.7% on GDAD, C-Eval, and TeleQnA, respectively; it exceeds LocMoE by 4.1%, 1.2%, and 1.7% on these datasets, respectively. Both theoretical proofs and experimental results demonstrate the benefits of the LocMoE+ routing strategy for LLM training efficiency. Future work may explore methods to compress communication data to further reduce communication overhead.
Appendix A Missing Proofs
A.1 Auxiurary Results
Lemma A.1 (Theorem 4 in [42]).
@@ A.1 Auxiurary Results
${\mathbb{P}}({\bm{X}}_{i}=1)=p_{i},{\mathbb{P}}({\bm{X}}_{i}=0)=1-p_{i}.$
We consider the sum ${\bm{X}}=\sum_{i=1}^{n}{\bm{X}}_{i}$, with expectation $\mathbb{E}({\bm{X}})=\sum_{i=1}^{n}p_{i}$. Then we have
$\mathrm{(Lower\ tail)}\quad{\mathbb{P}}({\bm{X}}\leq\mathbb{E}{\bm{X}}-\lambda)\leq e^{-\frac{\lambda^{2}}{2\mathbb{E}{\bm{X}}}},$ (9)
$\mathrm{(Upper\ tail)}\quad{\mathbb{P}}({\bm{X}}\geq\mathbb{E}{\bm{X}}+\lambda)\leq e^{-\frac{\lambda^{2}}{2(\mathbb{E}{\bm{X}}+\lambda/3)}}.$
A.2 Proof of Theorem 3.3
@@
1) For the TCR, denote $s_{i}=\left|\left\{t<k:{\bm{x}}_{t}\text{ sent to expert }i,{\bm{x}}_{k}={\bm{o}}_{i}\right\}\right|,\forall i\in[n]$ as the top class-irrelevant token number candidated to the $i$-th expert before the valid token. Then by Assumption 3.2, each class-irrelevant token uniformly gives to any expert, leading to $s_{i}|({\bm{x}}_{k}={\bm{o}}_{i})\sim{\mathcal{B}}(k-1,1/n)$ (Binomial distribution), that is,
${\mathbb{P}}(s_{i}=t|{\bm{x}}_{k}={\bm{o}}_{i})=\begin{pmatrix}k-1\\ t\end{pmatrix}\cdot\left(\frac{1}{n}\right)^{t}\left(1-\frac{1}{n}\right)^{k-1-t},\forall t\in[k-1].$
@@
$=\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}{\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(C+\sum_{k=C+1}^{s}{\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})\right).$ (10)
Note that $\mathbb{E}s_{i}=(k-1)/n$. When $k\geq 2nC$, by lower tail bound in Lemma A.1, we get
${\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})\leq e^{-\frac{(k-1-n(C-1))^{2}}{2(k-1)n}}\leq e^{-\frac{k-1}{8n}}.$
@@
$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\sum_{k=2nC+1}^{s}{\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})\right)$
$\leq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\sum_{k=2nC}^{s-1}e^{-\frac{k}{8n}}\right)\leq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\frac{e^{-\frac{C}{4}}}{1-e^{-\frac{1}{8n}}}\right)$
@@
Moreover, for $1+\frac{nC}{4}\leq k\leq 1+\frac{nC}{2}$, i.e., $2(k-1)\leq nC\leq 4(k-1)$, by upper tail bound in Lemma A.1, we get
${\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})=1-{\mathbb{P}}(s_{i}\geq C|{\bm{x}}_{k}={\bm{o}}_{i})\geq 1-e^{-\frac{3(nC-k+1)^{2}}{2n[2(k-1)+nC]}}\geq 1-e^{-\frac{k-1}{4n}}.$
@@
$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\sum_{k=\lceil 1+nC/4\rceil}^{\lfloor 1+nC/2\rfloor}{\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})\right)$
$\geq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\frac{nC}{4}-1-\sum_{k=\lceil 1+nC/4\rceil}^{\lfloor 1+nC/2\rfloor}e^{-\frac{k-1}{4n}}\right)\geq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\frac{nC}{4}-1-\frac{e^{-\frac{C}{16}}}{1-e^{-\frac{1}{4n}}}\right)$
@@
2) For the ECR, denote $s_{i}$ as the class-irrelevant token number with the score larger than ${\bm{o}}_{i}$ for $i$-th expert. By Assumption 3.2, we derive that $s_{i}\sim{\mathcal{B}}(s-1,q_{i}),\forall i\in[n]$.
@@
$=\frac{1}{n}\sum_{i=1}^{n}{\mathbb{P}}(s_{i}\leq C-1,s_{i}\sim{\mathcal{B}}(s-1,q_{i}))$
If $C-1\leq(s-1)q_{i}/2$, by lower tail bound in Lemma A.1 with $\lambda=(s-1)q_{i}-(C-1)<\mathbb{E}s_{i}$, we obtain that
${\mathbb{P}}(s_{i}\leq C-1)\leq e^{-\frac{(s-1)q_{i}}{2}\left(1-\frac{C-1}{(s-1)q_{i}}\right)^{2}}\leq e^{-\frac{(s-1)q_{i}}{8}}.$
If $C\geq 2(s-1)q_{i}$, by upper tail bound in Lemma A.1 with $\lambda=C-(s-1)q_{i}>0$, we obtain that
${\mathbb{P}}(s_{i}\leq C-1)=1-{\mathbb{P}}(s_{i}\geq C)\geq 1-e^{-\frac{[C-(s-1)q_{i}]^{2}}{2(C+2(s-1)q_{i})/3}}\geq 1-e^{-\frac{3C}{16}}.$

### v2

@@ Introduction
Hence, the traning of an input in the current step is valid if the class-discriminative patterns is correctly dispatched. To quantitatively measure the difference between TCR and ECR, we define training success rate of input motivated by the training process of MoE.
Definition 2 (training success rate).
We say the input $\bm{x}\in\mathcal{R}^{sd}$ with $s$ tokens succeed in training if the class-discriminative pattern in $\bm{x}$, e.g., $\bm{o}_{i}$ is correctly dispatched to $i$-th expert. We further define training success rate as the probability that the input succeed in training.
Furthermore, to show the quantitative comparison of TCR and ECR in training success rate, we need following asssumptions and notations of token patterns.
Assumption 3 (class-discriminative).
@@
$\mathcal{P}(\delta_{\bm{o}_{i},i}\geq\delta_{\bm{x}_{j},i},\forall j\in[s])=p_{i}\geq 1/n,\forall i\in[n].$
Assumption 4 (class-irrelevant).
@@
$\mathcal{P}(\bm{r}\sim\mathcal{N},\delta_{\bm{r},i}\geq\delta_{\bm{x}_{j},i},\forall j\in[s])=1/n,\forall i\in[n].$ (7)
@@
$\mathcal{P}(\bm{r}\sim\mathcal{N},\delta_{\bm{r},i}\geq\delta_{\bm{o}_{i},i})=q_{i},\forall i\in[n],$ (8)
@@
Assumption 3 assumes the valid token is uniformly distributed in training samples due to the massive amounts of data nowadays. Assumption 4 assumes the invalid tokens can be uniformly dispatched to experts since the invalid tokens do not provide supervised signal to router and experts in training. We consider such uniform settings are common assumptions in theoretical analyis. Now we compute the training success rate of TCR and ECR.
Theorem 5.
Under Assumptions 3 and 4, the training success rate of TCR in each sample $\bm{x}$ is
@@
and the training success rate of ECR is $\forall i\in[n]$,
@@
Corollary 6.
In practice, For constant number of experts (Jiang et al. 2024a), i.e., $n=\Theta(1)$, and $C<s$ to save computation cost. We have the following lower bound for capacity $C$ to ensure high training success rate:
  1. 1.
    Suppose $q_{i}=\Theta(1)$. Then TCR is much better than ECR, and we only need $C=\Theta(s)$.
  2. 2.
    Suppose $\forall i\in[n],sq_{i}\leq C^{*}$ for some $C^{*}>0$. Then ECR is much better than TCR, and we only need $C\geq 2C^{*}$.
Remark 7.
We explain the benefit of swicthing TCR to ECR during training based on Theorem 5 and feature distrution during training.
At the beginning of training, the model seldom learn the task. Then the feature of class-irrelevant tokens is nearly isotropy, e.g., uniformly distrbute around the sphere (see Appendix), leading to $q_{i}=\Theta(1)$. The succed rate of TCR with the form $C/s$ is better than ECR with the form $e^{-s}$. Thus we should choose TCR with a large capacity $C=\Theta(s)$ to improve the success rate of training samples.
After training for some iterations, the experts can roughly distinguish the class-irrelevant and discriminative patterns, leading to $q_{i}\ll 1$ or $sq_{i}\leq C^{*}$ for some $C^{*}>0$ (see Appendix). Then ECR with success rate nearly $1$ is better than TCR with the form $C/s$ as long as $C\geq 2C^{*}$. Thus we should choose ECR with a small capacity $C=\Theta(1)$ to improve the success rate of training samples.
Indeed, we find that Chowdhury et al. (2023, the definition of $\ell^{*}$) consider the ECR setting and verify the benefit in sample complexity. They assume the maximum number of class-irrelevant patches that are close to class-discriminative patches are bounded, which has similar effect as $C^{*}$ in our scene.
Communication Optimization
Experiments
Experimental Setup
This study employs the Mixtral 8×7B model, incorporating LocMoE and our proposed approach. The Mixtral model, comprising 46.7 billion parameters and utilizing Group Query Attention (GQA), features 32 sparse expert blocks with 8 experts in the MoE Feedforward layer, where each token engages the top 2 experts for processing. Given the prevalence of long-text corpora in our application scenarios, we extended the sequence length to 32,768 and implemented tailored parallel strategies for cluster scales of 32N, 64N, and 256N, encompassing tensor, pipeline, data, and expert parallelism, with a consistent global batch size of 128. Other details of experimental setup including datasets, environment, configuration, and metrics, can be seen in Appendix.
Efficiency Promotion and Memory Footprint Reduction
As detailed in Section Method, we consistently use Top-1 routing to ensure the routing implementation aligns with our theoretical framework. The Baseline model utilizes a limited expert capacity mode instead of the groupedGEMM scheme, which avoids token dropping, with the capacity factor set to 1.1. LocMoE considers data distribution uniformity and estimates expert capacity using a lower bound formula derived from its theoretical conclusions in the first batch, maintaining it as a constant during subsequent training. Our approach (abbreviate to ”LocMoE+” in figures) fixes the range of score sums, processes hidden states, and calculates current expert capacity. The subsequent analysis addresses the training time, convergence, and memory usage efficiency of these schemes on multiple sizes of Ascend clusters.
@@
Figure 3a illustrates the time consumption of these methods during the first 1000 iterations of training. Due to initialization and some unstable factors, time consumption is recorded starting from the 5th iteration. The Baseline model’s time consumption is relatively stable. As iterations increase, LocMoE’s time consumption slightly decreases, particularly in 32N and 64N, consistent with the conclusion that locality loss is effective only when the number of experts is greater than or equal to the number of nodes. Our approach incurs slightly higher time consumption than LocMoE due to the computational overhead of token rearrangement. However, as token features converge, the required tokens gradually decrease and stabilize, leading to a decline in time consumption, which remains stable in subsequent training processes. Overall, our approach reduces training time by 2.9% to 13.3% compared to LocMoE, and by 5.4% to 46.6% compared to the Baseline.
@@
We select 10 iterations at equal intervals from the training iterations to collect data on the time consumption of computation, communication, overlap, and idle periods, as shown in Figure 3(a). It is important to note that the data collection operation also introduces some overhead. After integrating LocMoE and our approach, the time consumption of each component decreases, with a significantly greater reduction in computation overhead compared to communication overhead. Additionally, as the cluster size increases, the proportion of computation/communication overlap decreases, and the magnitude of the reduction in computation overhead diminishes. Figure 3(b) illustrates perplexity as a measure of convergence. The convergence curves of these approaches indicate normal loss convergence, with our approach not adversely impacting convergence.
Figure 5: print recorded in one acquisition cycle with different schemes and cluster sizes.
@@
The proportional time consumption at the operator level is depicted in Figure 6. Among the components, AI CORE efficiently executes matrix multiplications and convolutions in AI algorithms; AI VECTOR CORE accelerates vector operations through parallel processing; MIX AIC integrates different types of operators and optimizes for multiple tasks; AI CPU is optimized in hardware and instruction sets to better support AI algorithms. Our approach selects fewer tokens, resulting in a 17$\times$ performance improvement in the FFN MatMul operator compared to the Baseline and a 2.6$\times$ improvement compared to LocMoE. This leads to an overall 2.8$\times$ reduction in the cumulative time consumption of the MatMul operator and a 2.6$\times$ decrease in Cube computing load. However, the proportions of TopK and IndexPutV2, related to rearrangement, show a slight increase.
@@
We select a single iteration during the stable training period and describe the per-device memory usage (Allocated) using the first 100,000 samples from its memory monitoring, as shown in Figure 5. Overall, our approach achieves memory usage reduction of 4.57% to 16.27% compared to the Baseline and 2.86% to 10.5% compared to LocMoE. As cluster size increases, the proportion of computational overhead decreases, and the gap in memory usage narrows. Additionally, instantaneous memory peaks gradually disappear, and the fluctuation amplitude of short-term memory also diminishes.
The Performance of Downstream Tasks
To enhance the model’s conversational capabilities and adaptability to downstream task, we fine-tuned the pre-trained models. As shown in Figure 7, with sufficient supervised fine-tuning (SFT), our approach achieves an average improvement of approximately 20.1% in 16 sub-capabilities of Domain Task Capability, which is a portion of General and Domain-specific Assessment Dataset (GDAD), compared to the Baseline, and an increase of about 3.5% compared to LocMoE. The Rewriting and Summary capabilities show the highest improvement, with a 28.2% increase compared to the Baseline and a 6.7% increase compared to LocMoE. In the 13 tests of Domain Competency Exam, our approach demonstrates an average improvement of 16% relative to the Baseline and an average increase of approximately 4.8% compared to LocMoE. The IP Training in the digital communications domain shows the most significant improvement, with a 27.3% increase compared to the Baseline and a 3.0% increase compared to LocMoE. Among the 18 sub-capabilities of General Ability, our approach exhibits an improvement of about 13.9% relative to the Baseline and an average increase of 4.8% compared to LocMoE. The capability of Planning demonstrates the highest improvement, with a 26.8% increase compared to the Baseline and a 2.92% increase compared to LocMoE.
Table 1 presents the holistic evaluation results for multiple datasets, where GDAD-1 represents Domain Task Capability, and the other metrics follow accordingly. Notably, due to the 6:4 ratio of Chinese to English data in our incremental pre-training domain data and the 7:3 ratio in the fine-tuning data, our approach achieves an improvement of approximately 10.7% compared to the Baseline and 1.2% compared to LocMoE in the C-Eval (Huang et al. 2024) evaluation, despite the limited data available for training. During incremental training and fine-tuning, we incorporated substantial telecommunications domain knowledge, questions, and case studies. TeleQnA (Maatouk et al. 2023), the first benchmark dataset designed to evaluate the knowledge of LLMs in telecommunications, effectively measures the model’s capabilities in this domain. Consequently, our approach comprehensively surpasses both the Baseline and LocMoE on this specific dataset.
Table 1: Performance promotion obtained by our approach on different datasets.
@@
Conclusion
In this paper, we propose a novel MoE structure incorporating token feature awareness mechanism to enhance the training efficiency and performance of LLMs. Building upon the feature extraction component and locality loss of LocMoE, our approach introduces a routing strategy that combines TCR and ECR, theoretically demonstrating improved training success rates. By defining the affinity between tokens and experts, expert capacity can be dynamically reduced and further minimized. Model training experiments are conducted on Ascend clusters, leveraging the communication optimization characteristics of CoC. Our approach achieves performance improvements up to 46.6% (32N) compared to the Baseline and 13.3% (32N) compared to LocMoE, while reducing memory usage by up to 16.27% and 10.5%, respectively. To evaluate model performance, all pre-trained models are fine-tuned using a mixture of domain-specific and general data, and assessed with the open-source datasets C-Eval and TeleQnA, and closed domain benchmark GDAD. In downstream tasks, our approach outperforms the Baseline by 14.1%, 10.7%, and 9.7% on GDAD, C-Eval, and TeleQnA, respectively; it exceeds LocMoE by 4.1%, 1.2%, and 1.7% on these datasets, respectively. Both theoretical proofs and experimental results demonstrate the benefits of the proposed routing strategy for LLM training efficiency. Future work may explore methods to compress communication data to further reduce communication overhead.
Appendix A Appendix
Appendix B Missing Proof
Auxiurary Results
Lemma 8 (Theorem 4 in (Chung and Lu 2006)).
@@ Appendix B Missing Proof
$\mathcal{P}(\bm{X}_{i}=1)=p_{i},\mathcal{P}(\bm{X}_{i}=0)=1-p_{i}.$
We consider the sum $\bm{X}=\sum_{i=1}^{n}X_{i}$, with expectation $\mathcal{E}(X)=\sum_{i=1}^{n}p_{i}$. Then we have
$\mathrm{(Lower\ tail)}\quad\mathcal{P}(\bm{X}\leq\mathcal{E}\bm{X}-\lambda)\leq e^{-\frac{\lambda^{2}}{2\mathcal{E}\bm{X}}},$ (11)
$\mathrm{(Upper\ tail)}\quad\mathcal{P}(\bm{X}\geq\mathcal{E}\bm{X}+\lambda)\leq e^{-\frac{\lambda^{2}}{2(\mathcal{E}\bm{X}+\lambda/3)}}.$
Proof of Theorem 5
@@
1) For the TCR, denote
$s_{i}=\left|\left\{t<k:\bm{x}_{t}\text{ sent to expert }i,\bm{x}_{k}=\bm{o}_{i}\right\}\right|,\forall i\in[n]$
as the top class-irrelevant token number candidated to the $i$-th expert before the valid token. Then by Assumption 4, each class-irrelevant token uniformly gives to any expert, leading to $s_{i}|(\bm{x}_{k}=\bm{o}_{i})\sim\mathcal{B}(k-1,1/n)$ (Binomial distribution), i.e., $\forall t\in[k-1]$,
$\mathcal{P}(s_{i}=t|\bm{x}_{k}=\bm{o}_{i})=\begin{pmatrix}k-1\\ t\end{pmatrix}\cdot\left(\frac{1}{n}\right)^{t}\left(1-\frac{1}{n}\right)^{k-1-t}.$
@@
$=\sum_{i=1}^{n}\mathcal{P}(\bm{o}_{i}\text{ sent to expert }i|\bm{o}_{i}\text{ is in }\bm{x})\cdot\mathcal{P}(\bm{o}_{i}\text{ is in }\bm{x})$
$=\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$
$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(C+\sum_{k=C+1}^{s}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right).$ (12)
Note that $\mathcal{E}s_{i}=(k-1)/n$. When $k\geq 2nC$, by lower tail bound in Lemma 8, we get
$\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\leq e^{-\frac{(k-1-n(C-1))^{2}}{2(k-1)n}}\leq e^{-\frac{k-1}{8n}}.$
@@
(12)${{\leq}}\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$ )
$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\sum_{k=2nC+1}^{s}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right)$
$\leq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\sum_{k=2nC}^{s-1}e^{-\frac{k}{8n}}\right)$
$\leq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\frac{e^{-\frac{C}{4}}}{1-e^{-\frac{1}{8n}}}\right)$
@@
Moreover, for $1+\frac{nC}{4}\leq k\leq 1+\frac{nC}{2}$, i.e., $2(k-1)\leq nC\leq 4(k-1)$, by upper tail bound in Lemma 8, we get
$\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$ $=1-\mathcal{P}(s_{i}\geq C|\bm{x}_{k}=\bm{o}_{i})$
$\geq 1-e^{-\frac{3(nC-k+1)^{2}}{2n[2(k-1)+nC]}}\geq 1-e^{-\frac{k-1}{4n}}.$
@@
(12)${{\geq}}\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$ )
$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\sum_{k=\lceil 1+nC/4\rceil}^{\lfloor 1+nC/2\rfloor}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right)$
$\geq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\frac{nC}{4}-1-\sum_{k=\lceil 1+nC/4\rceil}^{\lfloor 1+nC/2\rfloor}e^{-\frac{k-1}{4n}}\right)$
$\geq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\frac{nC}{4}-1-\frac{e^{-\frac{C}{16}}}{1-e^{-\frac{1}{4n}}}\right)$
@@
2) For the ECR, denote $s_{i}$ as the class-irrelevant token number with the score larger than $\bm{o}_{i}$ for $i$-th expert. By Assumption 4, we derive that $s_{i}\sim\mathcal{B}(s-1,q_{i}),\forall i\in[n]$.
@@
$=\sum_{i=1}^{n}\mathcal{P}(\text{expert }i\text{ choose }\bm{o}_{i}|\bm{o}_{i}\text{ is in }\bm{x})\mathcal{P}(\bm{o}_{i}\text{ is in }\bm{x})$
$=\frac{1}{n}\sum_{i=1}^{n}\mathcal{P}(s_{i}\leq C-1,s_{i}\sim\mathcal{B}(s-1,q_{i}))$
If $C-1\leq(s-1)q_{i}/2$, by lower tail bound in Lemma 8 with $\lambda=(s-1)q_{i}-(C-1)<\mathcal{E}s_{i}$, we obtain that
$\mathcal{P}(s_{i}\leq C-1)\leq e^{-\frac{(s-1)q_{i}}{2}\left(1-\frac{C-1}{(s-1)q_{i}}\right)^{2}}\leq e^{-\frac{(s-1)q_{i}}{8}}.$
If $C\geq 2(s-1)q_{i}$, by upper tail bound in Lemma 8 with $\lambda=C-(s-1)q_{i}>0$, we obtain that
$\mathcal{P}(s_{i}\leq C-1)$ $=1-\mathcal{P}(s_{i}\geq C)$
$\geq 1-e^{-\frac{[C-(s-1)q_{i}]^{2}}{2(C+2(s-1)q_{i})/3}}\geq 1-e^{-\frac{3C}{16}}.$